There are 3 piles of rocks, A, B and C where B has 2 more than A and C has 4 times as many rocks as A. The total number of rocks is 14. Find the number of rocks in each pile using x and to do the problem in 3 different ways, using the x in three different positions.

Answer by ankor@dixie-net.com(22740)

There are 3 piles of rocks, A, B and C where B has 2 more than A and C has 4 times as many rocks as A. The total number of rocks is 14. Find the number of rocks in each pile using x and to do the problem in 3 different ways, using the x in three different positions.
:
x = A pile
then
(x+2) = B pile
and
4x = C pile
:
x + (x+2) + 4x = 14
6x = 14 - 2
6x = 12
x = 2 rocks in A pile
:
You find the no. in B and C
:
Seems like a good way to do this